OpenStudy (anonymous):

HOW TO GRAPH: -3/2+ 2y = 8

OpenStudy (marco26):

You can write it first in the form y=mx+b where m is the slope and b is the y-intercept. Once donce, plot the y-intercept point, and from that point apply rise/run based on the slope m

OpenStudy (anonymous):

Since there is no x: move -3/2 to other side, 2y = 8 + 3/2 divide both sides by 2: y= 4 + 3/4 y= 19/4 graphic a horizontal line which passes y=19/4 (19/4 ~= 5)